what is eccentricity?


plzz explian with draw..........

Grade:11

1 Answers

vikas askiitian expert
13 years ago

Eccentricity is represented as a quality or amount of being strange or it’s a kind of state. The eccentricity is normally defined as the ratio of the distance between any points on the path on a conic part from where it’s threw and last point of the throw from the corresponding directrix. In general the eccentricity will look like a “throw”.

Description to Eccenctricity of an Ellipse:

The diagram will given below the structure of  an eccentricity of an ellipse:

eccentricity

from the diagram, we can know the formula for eccentricity, e=c/a

c=(a2 -b2)1/2

e=(a2 - b2)1/2 /a